OpenStudy (anonymous):

A rectangle has a perimeter of 26 feet. Twice its length (l) is one foot more than three times its width (w). What is the length of the rectangle, in feet? A. 5 B. 13 C. 8 D. 16 thank you to all who help me on this :)

OpenStudy (texaschic101):

P = 2L + 2W (formula for perimeter of rectangle) P = 26 2L = 3W + 1 L = (3W + 1)/2 now we sub 26 = 2(3W + 1)/2 + 2W 26 = 3W + 1 + 2W 26 = 5W + 1 26 - 1 = 5W 25 = 5W 5 = W 2L = 3W + 1 2L = 3(5) + 1 2L = 15 + 1 2L = 16 L = 8 check.. P = 2L + 2W 26 = 2(8) + 2(5) 26 = 16 + 10 26 = 26 (correct) Length is 8